# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 10.4 Strategic Buyers
Liu et al. (2024) introduce buyer manipulation into contextual dynamic pricing. At time $t$, a buyer arrives with true covariates $x_{t}^{0}\in\mathbb{R}^{d}$ and valuation $v_{t}=\theta_{0}^{\top}x_{t}^{0}+z_{t}$. The seller announces a pricing rule $p_{t}=g(\hat{\theta}_{k}^{\top}x_{t})$, where $\hat{\theta}_{k}$ is the current parameter estimate. Crucially, the buyer observes this rule and can distort her reported features. She solves a cost-minimization problem:
@@
where $A$ is a positive definite matrix governing manipulation costs.200200200The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
Theorem 6 (Theorem 1 of Liu et al. (2024)).
@@
The fix is to jointly estimate demand parameters and manipulation behavior. Liu et al. (2024) propose an episodic algorithm with two phases per episode. During the exploration phase, the seller posts uniform random prices that do not depend on features. Since the price is independent of $\tilde{x}_{t}$, buyers have no incentive to manipulate, and the seller observes true features $x_{t}^{0}$. During the exploitation phase, the seller uses the corrected pricing rule that anticipates the manipulation:
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.201201201Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 10.5 Comparison of Regret Rates
Liu et al. (2024) linear ++ strategic known, naïve $T$ never improves
Liu et al. (2024) linear ++ strategic known, corrected $d\sqrt{T}$ $\sim$500 lost
@@ 10.6.1 Joint Assortment and Pricing at Scale
Ganti et al. (2018) deploy Thompson Sampling for dynamic pricing at Walmart.com. Their MAX-REV-TS algorithm models demand for each item $i$ on day $t$ via a constant-elasticity function $d_{i,t}(p)=f_{i,t}(p/p_{i,t-1})^{\gamma_{i}^{*}}$, where $d_{i,t}(p)$ is unit sales at price $p$, $f_{i,t}$ is a baseline demand forecast at the previous day’s price $p_{i,t-1}$, and $\gamma_{i}^{*}<-1$ is the unknown price elasticity. The structural assumption, that demand responds to price through a single elasticity parameter per item, reduces the learning problem from estimating a full demand curve to estimating one scalar per item. MAX-REV-TS places a Gaussian prior over the elasticity vector $\gamma^{*}$ and draws posterior samples at each period to solve a constrained revenue maximization problem. In a five-week field experiment on a basket of roughly 5,000 items, Thompson Sampling produced a statistically significant increase in per-item revenue relative to the passive pricing baseline.
@@ 10.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.202202202The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates203203203At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@ 11 Offline Reinforcement Learning and Human Feedback
This is the problem of offline reinforcement learning, also called batch reinforcement learning.204204204The term “batch RL” was standard in the earlier literature (Ernst et al., 2005; Lange et al., 2012). “Offline RL” became dominant after Levine et al. (2020), who distinguished it from off-policy RL (which still collects new data, just under a different policy than the target). I use “offline RL” throughout. The agent never queries the environment. All learning happens from a fixed dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})\}_{i=1}^{n}$ collected by some behavioral policy $\pi_{b}$. The goal is to find a policy $\hat{\pi}$ whose value $V^{\hat{\pi}}$ is as close to the optimal $V^{*}$ as possible.
Online RL algorithms (Q-learning, SARSA, policy gradient methods from Section 4) can in principle be applied to offline data by treating the dataset as a replay buffer. In practice, this fails catastrophically. The reason is distributional shift: the learned policy $\hat{\pi}$ inevitably queries state-action pairs that the behavioral policy $\pi_{b}$ never visited, and the Q-function at these unseen pairs is pure extrapolation. Because the Bellman backup propagates these extrapolation errors through the max operator, errors compound geometrically across the planning horizon, producing arbitrarily poor policies.205205205This failure mode was first demonstrated empirically by Fujimoto et al. (2019), who showed that standard off-policy algorithms (DDPG, SAC) trained purely from a static dataset performed worse than the behavioral policy itself, even when that behavioral policy was a partially-trained, mediocre agent. The gap widened with dataset size, the opposite of what one expects from more data.
@@ 11.2.1 Fitted Q-Iteration
Fitted Q-Iteration (FQI, Ernst et al., 2005) is the simplest offline RL algorithm, predating the modern pessimism framework. FQI applies the standard Bellman backup iteratively using supervised regression on the fixed dataset, as described in Section 4. It does not include any explicit pessimism mechanism. When state-action coverage is poor, the max in the Bellman backup selects the highest Q-value among all actions at $s^{\prime}$, including actions never observed in the data. If the function approximator generalizes poorly at these unseen actions, targets become noisy and biased upward, causing the overestimation cascade. FQI works well when coverage is good but degrades as coverage gaps grow.206206206Munos and Szepesvári (2008) prove finite-time error bounds for FQI under approximate Bellman completeness and all-policy concentrability. Both assumptions are strong, and violation of either leads to divergence in practice.
@@ 11.3 Simulation Study: Offline RL for Dynamic Pricing
The simulation study evaluates the four offline RL algorithms on a perishable inventory pricing problem with demand regime switching. A retailer with $I_{\max}=30$ units of perishable inventory must set prices over $H=20$ periods. The state $(i,d,t)$ consists of current inventory $i\in\{0,\ldots,30\}$, demand regime $d\in\{1,2,3,4\}$, and time remaining $t\in\{1,\ldots,20\}$. The action is a price $p\in\{1,\ldots,10\}$. Demand follows $Q\sim\text{Poisson}(\lambda_{0}[d]\cdot e^{-0.15p})$ with base rates $\lambda_{0}=(1.5,3.0,5.0,8.0)$, and the reward is $r=p\cdot\min(Q,i)$. Demand regimes follow a 4-state Markov chain with diagonal persistence 0.6. Unsold inventory at the terminal period incurs a spoilage cost of \$2.00 per unit, making clearance pricing valuable near the deadline.207207207The spoilage penalty creates distributional shift. The optimal policy adapts prices to inventory level and time remaining, using lower prices near the deadline when inventory is high. The behavioral policy ignores these state variables and prices at the maximum, so the Q-function at state-adapted pricing actions is extrapolation from sparse data. The \$2.00 penalty is a deliberate design choice: under harsher penalties (e.g., \$10 per unit), all methods collapse to 48–53% of optimal regardless of algorithmic sophistication, confirming that no offline correction can overcome severe distributional shift when the penalty regime amplifies consequences of the behavioral policy’s suboptimality. The behavioral policy represents a conservative pricing team that always sets the maximum price ($p=10$) regardless of demand regime, inventory, or time remaining, with probability 0.85 and randomizes uniformly over all prices with probability 0.15.208208208With 500 episodes (10,000 transitions) on a state-action space of 24,800 pairs, the 85% concentration at price 10 ensures that the behavioral state-action occupancy diverges significantly from the optimal policy’s occupancy, while the 15% uniform component provides sparse off-policy coverage. All episodes start at full inventory ($i=30$). All offline methods train on 500 episodes and are evaluated over 1,000 episodes against the DP optimal policy computed by backward induction.209209209FQI uses the standard Bellman backup with $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$, deliberately without a target network, to isolate the overestimation cascade as a pedagogical baseline. Adding target networks to FQI mitigates but does not eliminate extrapolation error. CQL and IQL include target networks following their original implementations (Kumar et al., 2020; Kostrikov et al., 2022); for CQL, target networks proved essential, as the conservative penalty amplifies bootstrap instability without them. Results are averaged over 20 independent seeds.
@@
FQI achieves 81.2% of the DP optimal, substantially below the behavioral cloning baseline of 88.0% (Table 20). Without any mechanism to control extrapolation error, the $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ operator in FQI’s Bellman backup selects overestimated Q-values at out-of-distribution actions, and these overestimates compound across 200 iterations of fitted Q-iteration. CQL and IQL both exceed the behavioral baseline, achieving 91.9% and 92.0% respectively.210210210CQL uses $\alpha=0.1$, the result of a search over $\alpha\in\{5.0,2.0,0.5,0.1\}$. Larger values push Q-values down too aggressively, collapsing the learned policy to the behavioral action at most states; the right $\alpha$ is problem-specific and can vary by orders of magnitude. CQL’s conservative penalty suppresses Q-values at actions not well-represented in the data while preserving relative ordering among data-supported actions, allowing the policy to discover lower prices that clear inventory near the deadline. IQL achieves the same improvement through a different mechanism: by replacing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ with the expectile-regressed value function $V(s^{\prime})$ as the Bellman continuation, IQL avoids querying Q at unseen actions during training while still extracting a policy that improves on the behavioral at states where the 15% noise component revealed better pricing actions. BCQ matches the behavioral at 88.0%; its action constraint restricts the policy to prices near 10, preventing both overestimation and improvement.211211211When the behavioral policy concentrates 85% probability at a single action, BCQ’s threshold constraint permits only that action at most states, effectively reducing BCQ to behavioral cloning regardless of the learned Q-values.
@@
Figure 21 varies the behavioral noise parameter $\epsilon_{b}\in\{0.05,0.3,0.9\}$. CQL and IQL remain above the BC baseline across all coverage levels, confirming that both pessimism mechanisms provide robustness to the data distribution. FQI exhibits non-monotone behavior: it peaks at moderate coverage ($\epsilon_{b}=0.3$) where partial exploration controls the overestimation cascade, but collapses at both extremes.212212212At high $\epsilon_{b}$, the near-uniform behavioral policy provides Q-function targets across all actions, giving the unconstrained $\max_{a^{\prime}}$ operator more opportunities to select overestimated values rather than fewer. BCQ collapses at $\epsilon_{b}=0.9$ because a nearly uniform behavioral policy renders the action constraint vacuous, reducing BCQ to unconstrained FQI. These coverage patterns connect directly to the concentrability framework (Definition D2): as $\epsilon_{b}$ decreases, the concentrability coefficient $C^{*}$ grows because the optimal policy’s state-action occupancy diverges from the behavioral policy’s, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (103) become large at states the optimal policy visits. CQL and IQL remain robust because their conservative adjustments scale with data sparsity, while FQI lacks this adaptive correction.

### 24 Mar 2026 v5

@@ 10.4 Strategic Buyers
Liu et al. (2024a) introduce buyer manipulation into contextual dynamic pricing. At time $t$, a buyer arrives with true covariates $x_{t}^{0}\in\mathbb{R}^{d}$ and valuation $v_{t}=\theta_{0}^{\top}x_{t}^{0}+z_{t}$. The seller announces a pricing rule $p_{t}=g(\hat{\theta}_{k}^{\top}x_{t})$, where $\hat{\theta}_{k}$ is the current parameter estimate. Crucially, the buyer observes this rule and can distort her reported features. She solves a cost-minimization problem:
@@
where $A$ is a positive definite matrix governing manipulation costs.202202202The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
Theorem 6 (Theorem 1 of Liu et al. (2024a)).
@@
The fix is to jointly estimate demand parameters and manipulation behavior. Liu et al. (2024a) propose an episodic algorithm with two phases per episode. During the exploration phase, the seller posts uniform random prices that do not depend on features. Since the price is independent of $\tilde{x}_{t}$, buyers have no incentive to manipulate, and the seller observes true features $x_{t}^{0}$. During the exploitation phase, the seller uses the corrected pricing rule that anticipates the manipulation:
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.203203203Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 10.5 Comparison of Regret Rates
Liu et al. (2024a) linear ++ strategic known, naïve $T$ never improves
Liu et al. (2024a) linear ++ strategic known, corrected $d\sqrt{T}$ $\sim$500 lost
@@ 10.6.1 Joint Assortment and Pricing at Scale
Ganti et al. (2018) deploy Thompson Sampling in-field for dynamic pricing at Walmart.com. Their MAX-REV-TS algorithm models demand for each item $i$ on day $t$ via a constant-elasticity function $d_{i,t}(p)=f_{i,t}(p/p_{i,t-1})^{\gamma_{i}^{*}}$, where $d_{i,t}(p)$ is unit sales at price $p$, $f_{i,t}$ is a baseline demand forecast at the previous day’s price $p_{i,t-1}$, and $\gamma_{i}^{*}<-1$ is the unknown price elasticity. The structural assumption, that demand responds to price through a single elasticity parameter per item, reduces the learning problem from estimating a full demand curve to estimating one scalar per item. MAX-REV-TS places a Gaussian prior over the elasticity vector $\gamma^{*}$ and draws posterior samples at each period to solve a constrained revenue maximization problem. In a five-week in-field experiment on a basket of roughly 5,000 items, Thompson Sampling produced a statistically significant increase in per-item revenue relative to the passive pricing baseline.
@@ 10.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.204204204The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates205205205At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
@@ 11 Offline Reinforcement Learning and Human Feedback
This is the problem of offline reinforcement learning, also called batch reinforcement learning.206206206The term “batch RL” was standard in the earlier literature (Ernst et al., 2005; Lange et al., 2012). “Offline RL” became dominant after Levine et al. (2020), who distinguished it from off-policy RL (which still collects new data, just under a different policy than the target). I use “offline RL” throughout. The agent never queries the environment. All learning happens from a fixed dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i},s_{i}^{\prime})\}_{i=1}^{n}$ collected by some behavioral policy $\pi_{b}$. The goal is to find a policy $\hat{\pi}$ whose value $V^{\hat{\pi}}$ is as close to the optimal $V^{*}$ as possible.
Online RL algorithms (Q-learning, SARSA, policy gradient methods from Section 4) can in principle be applied to offline data by treating the dataset as a replay buffer. In practice, this fails catastrophically. The reason is distributional shift: the learned policy $\hat{\pi}$ inevitably queries state-action pairs that the behavioral policy $\pi_{b}$ never visited, and the Q-function at these unseen pairs is pure extrapolation. Because the Bellman backup propagates these extrapolation errors through the max operator, errors compound geometrically across the planning horizon, producing arbitrarily poor policies.207207207This failure mode was first demonstrated empirically by Fujimoto et al. (2019), who showed that standard off-policy algorithms (DDPG, SAC) trained purely from a static dataset performed worse than the behavioral policy itself, even when that behavioral policy was a partially-trained, mediocre agent. The gap widened with dataset size, the opposite of what one expects from more data.
@@ 11.2.1 Fitted Q-Iteration
Fitted Q-Iteration (FQI, Ernst et al., 2005) is the simplest offline RL algorithm, predating the modern pessimism framework. FQI applies the standard Bellman backup iteratively using supervised regression on the fixed dataset, as described in Section 4. It does not include any explicit pessimism mechanism. When state-action coverage is poor, the max in the Bellman backup selects the highest Q-value among all actions at $s^{\prime}$, including actions never observed in the data. If the function approximator generalizes poorly at these unseen actions, targets become noisy and biased upward, causing the overestimation cascade. FQI works well when coverage is good but degrades as coverage gaps grow.208208208Munos and Szepesvári (2008) prove finite-time error bounds for FQI under approximate Bellman completeness and all-policy concentrability. Both assumptions are strong, and violation of either leads to divergence in practice.
@@ 11.3 Simulation Study: Offline RL for Dynamic Pricing
The simulation study evaluates the four offline RL algorithms on a perishable inventory pricing problem with demand regime switching. A retailer with $I_{\max}=30$ units of perishable inventory must set prices over $H=20$ periods. The state $(i,d,t)$ consists of current inventory $i\in\{0,\ldots,30\}$, demand regime $d\in\{1,2,3,4\}$, and time remaining $t\in\{1,\ldots,20\}$. The action is a price $p\in\{1,\ldots,10\}$. Demand follows $Q\sim\text{Poisson}(\lambda_{0}[d]\cdot e^{-0.15p})$ with base rates $\lambda_{0}=(1.5,3.0,5.0,8.0)$, and the reward is $r=p\cdot\min(Q,i)$. Demand regimes follow a 4-state Markov chain with diagonal persistence 0.6. Unsold inventory at the terminal period incurs a spoilage cost of \$2.00 per unit, making clearance pricing valuable near the deadline.209209209The spoilage penalty creates distributional shift. The optimal policy adapts prices to inventory level and time remaining, using lower prices near the deadline when inventory is high. The behavioral policy ignores these state variables and prices at the maximum, so the Q-function at state-adapted pricing actions is extrapolation from sparse data. The \$2.00 penalty is a deliberate design choice: under harsher penalties (e.g., \$10 per unit), all methods collapse to 48–53% of optimal regardless of algorithmic sophistication, confirming that no offline correction can overcome severe distributional shift when the penalty regime amplifies consequences of the behavioral policy’s suboptimality. The behavioral policy represents a conservative pricing team that always sets the maximum price ($p=10$) regardless of demand regime, inventory, or time remaining, with probability 0.85 and randomizes uniformly over all prices with probability 0.15.210210210With 500 episodes (10,000 transitions) on a state-action space of 24,800 pairs, the 85% concentration at price 10 ensures that the behavioral state-action occupancy diverges significantly from the optimal policy’s occupancy, while the 15% uniform component provides sparse off-policy coverage. All episodes start at full inventory ($i=30$). All offline methods train on 500 episodes and are evaluated over 1,000 episodes against the DP optimal policy computed by backward induction.211211211FQI uses the standard Bellman backup with $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$, deliberately without a target network, to isolate the overestimation cascade as a pedagogical baseline. Adding target networks to FQI mitigates but does not eliminate extrapolation error. CQL and IQL include target networks following their original implementations (Kumar et al., 2020; Kostrikov et al., 2022); for CQL, target networks proved essential, as the conservative penalty amplifies bootstrap instability without them. Results are averaged over 20 independent seeds.
@@
FQI achieves 81.2% of the DP optimal, substantially below the behavioral cloning baseline of 88.0% (Table 20). Without any mechanism to control extrapolation error, the $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ operator in FQI’s Bellman backup selects overestimated Q-values at out-of-distribution actions, and these overestimates compound across 200 iterations of fitted Q-iteration. CQL and IQL both exceed the behavioral baseline, achieving 91.9% and 92.0% respectively.212212212CQL uses $\alpha=0.1$, the result of a search over $\alpha\in\{5.0,2.0,0.5,0.1\}$. Larger values push Q-values down too aggressively, collapsing the learned policy to the behavioral action at most states; the right $\alpha$ is problem-specific and can vary by orders of magnitude. CQL’s conservative penalty suppresses Q-values at actions not well-represented in the data while preserving relative ordering among data-supported actions, allowing the policy to discover lower prices that clear inventory near the deadline. IQL achieves the same improvement through a different mechanism: by replacing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ with the expectile-regressed value function $V(s^{\prime})$ as the Bellman continuation, IQL avoids querying Q at unseen actions during training while still extracting a policy that improves on the behavioral at states where the 15% noise component revealed better pricing actions. BCQ matches the behavioral at 88.0%; its action constraint restricts the policy to prices near 10, preventing both overestimation and improvement.213213213When the behavioral policy concentrates 85% probability at a single action, BCQ’s threshold constraint permits only that action at most states, effectively reducing BCQ to behavioral cloning regardless of the learned Q-values.
@@
Figure 21 varies the behavioral noise parameter $\epsilon_{b}\in\{0.05,0.3,0.9\}$. CQL and IQL remain above the BC baseline across all coverage levels, confirming that both pessimism mechanisms provide robustness to the data distribution. FQI exhibits non-monotone behavior: it peaks at moderate coverage ($\epsilon_{b}=0.3$) where partial exploration controls the overestimation cascade, but collapses at both extremes.214214214At high $\epsilon_{b}$, the near-uniform behavioral policy provides Q-function targets across all actions, giving the unconstrained $\max_{a^{\prime}}$ operator more opportunities to select overestimated values rather than fewer. BCQ collapses at $\epsilon_{b}=0.9$ because a nearly uniform behavioral policy renders the action constraint vacuous, reducing BCQ to unconstrained FQI. These coverage patterns connect directly to the concentrability framework (Definition D2): as $\epsilon_{b}$ decreases, the concentrability coefficient $C^{*}$ grows because the optimal policy’s state-action occupancy diverges from the behavioral policy’s, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (103) become large at states the optimal policy visits. CQL and IQL remain robust because their conservative adjustments scale with data sparsity, while FQI lacks this adaptive correction.